Let x = age in years of a rural Quebec woman at the time of her first marriage. In the year 1941, the population variance of x was approximately σ2 = 5.1. Suppose a recent study of age at first marriage for a random sample of 41 women in rural Quebec gave a sample variance s2 = 2.7. Use a 5% level of significance to test the claim that the current variance is less than 5.1.

(a) What is the level of significance?


State the null and alternate hypotheses.

H0: σ2 = 5.1; H1: σ2 > 5.1H0: σ2 = 5.1; H1: σ2 < 5.1    H0: σ2 < 5.1; H1: σ2 = 5.1H0: σ2 = 5.1; H1: σ2 ≠ 5.1


(b) Find the value of the chi-square statistic for the sample. (Round your answer to two decimal places.)


What are the degrees of freedom?


What assumptions are you making about the original distribution?

We assume a normal population distribution.We assume a exponential population distribution.    We assume a binomial population distribution.We assume a uniform population distribution.


(c) Find or estimate the P-value of the sample test statistic.

P-value > 0.1000.050 < P-value < 0.100    0.025 < P-value < 0.0500.010 < P-value < 0.0250.005 < P-value < 0.010P-value < 0.005


(d) Based on your answers in parts (a) to (c), will you reject or fail to reject the null hypothesis?

Since the P-value > α, we fail to reject the null hypothesis.Since the P-value > α, we reject the null hypothesis.    Since the P-value ≤ α, we reject the null hypothesis.Since the P-value ≤ α, we fail to reject the null hypothesis.


(e) Interpret your conclusion in the context of the application.

At the 5% level of significance, there is insufficient evidence to conclude that the variance of age at first marriage is less than 5.1.At the 5% level of significance, there is sufficient evidence to conclude that the variance of age at first marriage is less than 5.1.  
Answer #1

a)

level of significance=0.05

H0: σ2 = 5.1; H1: σ2 < 5.1

b)

value of the chi-square statistic =(n-1)*(s/σ)2 =(41-1)*(2.7/5.1)=21.18

degrees of freedom =n-1=41-1=40

We assume a normal population distribution.

c)0.005 <P-value < 0.010

d)

since the P-value ≤ α, we reject the null hypothesis

e)

At the 5% level of significance, there is sufficient evidence to conclude that the variance of age at first marriage is less than 5.1.